OpenStudy (anonymous):

Extracting the roots a^2+12a+36=49

OpenStudy (anonymous):

\[(a+6)^2=49\] \[a+6=7\] so \[a=1\] OR \[a+6=-7\]so \[a=-13\]

OpenStudy (anonymous):

trick was to recognize that \[a^2+12a+36=(a+6)^2\] and you see it because \[a^2\] is the square of \[a\] 6 is the square of 36 and 12a is 2 times 6 times a

OpenStudy (anonymous):

Of course, using the general formula for solving quadratic equations works as well, but recognizing the square of a binomial is indeed a handy way to solve some equations faster. :)

OpenStudy (anonymous):

It is actually much easier to move the 49 over first \[a^2 + 12a - 13 = 0\] \[(a+13)(a-1)=0\]

OpenStudy (anonymous):

JethroKuan: I think it depends _a lot_ on the specific curriculum and school how much people practise factorising. I can see how you did that but by no means would have I did that myself. :) That method takes some practice of routines and tricks.